### v1

# Dark Matter from Holography

Abstract
Previous studies have examined the holographic principle as a means of producing dark energy. Here we propose instead the possibility of holographic dark matter. In this case, dark matter does not arise in the framework of particle physics but is derived from the infrared cutoff set by the horizon scale. Using the Ricci cutoff, and a universe containing only baryons and radiation, we can account for the dark matter and naturally explain the coincidence between baryonic and nonbaryonic contributions to the density. In the presence of a pre-existing vacuum energy density our model reverses the sign of this density, thus accounting for the fact that certain string theories generically predict a negative vacuum energy, but observations require a positive value.
preprint: APS/123-QED
Observations over the past few decades have yielded a standard model of cosmology, in which the universe has a flat geometry with $\Omega=1$ and consists of roughly 5% baryonic matter, 25% nonbaryonic dark matter [1, 2, 3], and 70% dark energy [4, 5]. While this picture shows remarkably agreement with observations, it leaves open the question of the nature of the nonbaryonic dark matter and the dark energy. (Henceforth we will use “dark matter” to refer only to nonbaryonic dark matter, eliding the fact that a significant fraction of the baryonic matter is also dark). A major effort has gone into the quest to discover the dark matter, including accelerator-based detection attempts, direct underground searches, and indirect searches, but no candidates have yet been detected [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16]. As various windows in parameter space have closed, it seems opportune to explore more exotic possibilities.
In this paper, we consider the possibility that dark matter arises as a corollary of the holographic principle [17, 18, 19]. The idea of a holographic energy density in the cosmological context has long been applied to dark energy. Cohen et al. [20] formulated the holographic energy bound by arguing that an effective quantum field theory with an ultraviolet cutoff $\Lambda$ defined in a region of size $L$ should not contain states that would collapse to form a black hole. This leads to the constraint that the total vacuum energy in the volume $L^{3}$ should not exceed the mass of a black hole with the same size. In terms of the vacuum energy $\rho_{\Lambda}$, this requirement takes the form
$L^{3}\rho_{\Lambda}\leq m_{Pl}^{2}L,$ (1)
where $m_{Pl}$ is the Planck mass. If this inequality is saturated, one obtains the relation for the holographic dark energy density, $\rho_{HDE}$:
$\rho_{HDE}=3c^{2}m_{Pl}^{2}L^{-2},$ (2)
where $c$ is a dimensionless constant of order unity that accounts for uncertainties in the saturation of the bound. This has led to huge body of literature devoted to studying these holographic models [21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31]. The reader should also note that in the analysis henceforth, we will set $m_{Pl}=1$.
Now consider the possibility that Eq. (2) gives not the dark energy density, but the dark matter density, which we will call holographic dark matter (HDM). The physical picture of HDM is that the dark matter density should not be thought of as an independent, adjustable component, but rather something determined through a horizon-based relation akin to the holographic inequality. Later we will determine whether both the dark energy and dark matter can be incorporated into a unified holographic picture.
Our starting point will be Eq. (2), rewritten to describe the HDM:
$\rho_{HDM}=3c^{2}L^{-2}.$ (3)
The simplest (and first proposed) choice for $L$ is the inverse Hubble distance,
$L=H^{-1},$ (4)
where $H=\dot{a}/a$, and $a$ is the scale factor. However, this cutoff will not be suitable for our purposes. Instead, we will make use of the Granda-Oliveros cutoff, given by [23]
$L=\left(\alpha H^{2}+\beta\dot{H}\right)^{-1/2},$ (5)
where $\alpha$ and $\beta$ are constants of $\mathcal{O}(1)$.
The holographic energy density given by Eqs. (3) and (5) will be dependent on the other components driving the Hubble expansion. We will take a minimal approach here; the only components we can absolutely account for with standard physics are the baryon density $\rho_{B}$ and the background radiation density $\rho_{R}$, which consists of both photons and neutrinos. With those assumptions, the Friedman equation becomes
$3H^{2}=3c^{2}(\alpha H^{2}+\beta\dot{H})+\rho_{B0}a^{-3}+\rho_{R0}a^{-4},$ (6)
where the 0 subscript denotes present-day values and we normalize the scale factor to $a=1$ today. We have set $c=1$ here, since it is of order unity and can in any case be absorbed into the definitions of $\alpha$ and $\beta$. Equation (6) integrates to
$H^{2}$ $=$ $\frac{1}{3-3\alpha+(9/2)\beta}\rho_{B0}a^{-3}+\frac{1}{3-3\alpha+6\beta}\rho_{R0}a^{-4}$ (7)
$+$ $Ka^{2(1-\alpha)/\beta},$
where $K$ is a constant of integration. Then our expression for $\rho_{HDM}$ from Eqs. (3) and (5) gives
$\rho_{HDM}$ $=$ $\frac{2\alpha-3\beta}{2-2\alpha+3\beta}\rho_{B0}a^{-3}+\frac{\alpha-2\beta}{1-\alpha+2\beta}\rho_{R0}a^{-4}$ (8)
$+$ $3Ka^{2(1-\alpha)/\beta},$
which was first derived in Ref. [23]. Thus, the Granda-Oliveros cutoff leads to additional components with radiation-like and matter-like evolution, respectively.
However, there are severe observational constants on any additional component that scales as radiation. [Note that Ref.[32] proposed a model with the Granda-Oliveros cutoff as a unified model for dark matter and dark energy, but such a model will inevitably violate the bounds on additional radiation]. While a small additional radiation component can be accommodated [33], it is simplest to take $\beta=\alpha/2$, yielding the holographic Ricci model, which was first proposed by Gao et al. [34] and independently by Cai et al. [35]. In this case, the expression for $\rho_{HDM}$ becomes
$\rho_{HDM}=\frac{\alpha}{4-\alpha}\rho_{B0}a^{-3}+3Ka^{4(1-\alpha)/\alpha}.$ (9)
The holographic model with this Ricci cutoff has been extensively studied as a model for dark energy [36, 37, 38, 39, 32]; in this case, the second term in Eq. (9), with an appropriate choice for $\alpha$, serves as the dark energy.
Here, however, we consider a different model, in which it is the dark matter, rather than the dark energy, which arises from the Ricci cutoff. In that case we fix the value of $\alpha$ so that the first term in Eq. (9) gives the correct value for $\rho_{DM}$. From Ref. [33], we have $\rho_{DM}/\rho_{b}=5.3-5.4$, which can be achieved by taking
$\alpha\approx 3.3-3.4.$ (10)
Since $\alpha$ in this model is expected to be of order unity, we see that the holographic dark matter model provides a natural explanation for the dark matter-baryon coincidence, as it naturally predicts their densities to have the same order of magnitude.
What of the second term in Eq. (9)? The parameter $K$ enters as a constant of integration and so is free to have any value. With the value of $\alpha$ fixed to provide the correct dark matter density, this second energy density decreases slightly more slowly than matter. As there is no observational evidence for such an additional component, we can simply set $K$ sufficiently small to evade any observational constraints. It goes without saying that the future observation of a small component with this evolution would represent a smoking gun for this model, although it would be an extraordinary coincidence for this energy density to currently lie just below the threshold of detection. We assume it to be more likely that the density of this component is negligible or zero.
Of course, it is necessary but not sufficient for the density of any dark matter component to scale as $a^{-3}$; it must also cluster gravitationally. A number of investigators have examined perturbation growth in holographic dark energy models [40, 39, 41]. While it is clear that the adiabatic sound speed, $c_{a}$, in our model is zero, that is not the sound speed $c_{s}$ that determines the clustering properties of the HDM. As emphasized in Ref. [41], it is impossible to determine $c_{s}$ without an underlying Lagrangian for the theory. In the absence of such a Lagrangian, previous investigators have generally inserted a value of $c_{s}$ by hand. We simply note here that $c_{s}$ in our model must be zero in order for the HDM to behave as dark matter.
If Ricci holographic dark matter serves as the dark matter alone, then we are left to assume a different source for the dark energy. However, any dark energy inserted into the Friedman equation will necessary receive an additional contribution from the holographic cutoff in Eq. (6). Consider the simplest case, where we have a preexisting “bare” cosmological constant with constant density $\rho_{\Lambda}$. Inserting this into Eq. (6) and solving for $\rho_{HDM}$ transforms Eq. (9) into
$\rho_{HDM}=\frac{\alpha}{4-\alpha}\rho_{B0}a^{-3}+3Ka^{4(1-\alpha)/\alpha}+\frac{\alpha}{1-\alpha}\rho_{\Lambda}.$ (11)
Adding together the bare cosmological constant and the holographic contribution from Eq. (11), we obtain a total effective cosmological constant given by
$\rho_{\Lambda eff}=\frac{1}{1-\alpha}\rho_{\Lambda}\approx-0.4\rho_{\Lambda}.$ (12)
Thus, in order for $\rho_{\Lambda eff}$ to be consistent with observations, the bare $\rho_{\Lambda}$ must be negative. This is actually a desirable feature of our model, as string theories generically predict a negative cosmological constant, and considerable effort has gone into converting this into the positive value of $\rho_{\Lambda}$ that we actually observe [42, 43, 44, 45, 46].
Note that it is possible, in principle, to derive holographic dark matter from other forms for the cutoff. For example, we could use the Hubble cutoff from Eq. (4). However, in this case we would again run into the issue of an additional radiation component. This can be evaded by taking the value of $c$ that appears in Eq. (2) to be a function of time, but it is clear that such models are considerably more contrived than the simple Ricci cutoff.
The model we have presented here provides a novel mechanism to generate dark matter, and it leads to the generation of a positive effective cosmological constant from a preexisting negative $\rho_{\Lambda}$. Further, it provides a natural mechanism for the rough equivalence of the baryon and dark matter densities. The key unresolved issue is whether such a form for dark matter can cluster as observed. A resolution of this question would require the development of an underlying Lagrangian for the holographic model.
The idea of holographic dark matter represents a radical paradigm shift. Instead of being composed of constituent particles, remnants, or extended objects, dark matter emerges as an information-theoretic fluid rooted in the holographic structure of spacetime itself. In particle-based frameworks, the dark matter density is determined by microphysical parameters such as mass, cross-sections, and freeze-out histories whereas in the holographic framework the density arises from the information capacity of the cosmic horizon. In this view, the dark matter content of the Universe is a macroscopic manifestation of the maximum allowable information encoded on cosmological horizons. In this perspective, we reinterpret the dark sector as an emergent phenomenon of the Universe’s boundary degrees of freedom rather than a sum of microscopic constituents in its bulk. To make this information-theoretic picture explicit one may recall that in quantum gravity the number of available microstates in any gravitational system scales as
$\mathcal{N}(L)\sim\exp\left[\frac{A(L)}{4}\right]$ (13)
where $A(L)$ is the area of the horizon measured in Planck units, and when this entropy is converted into an effective energy density via the standard relation $E\sim L^{-1}\mathcal{N}(L)$ one obtains the characteristic holographic scaling
$\rho_{\rm holo}(L)\propto L^{-2}$ (14)
which is the same structural dependence that underlies the HDM construction in Eq. (3). Thus, HDM can be viewed physically as the bulk gravitational response to the finite number of microstates allowed by the cosmic horizon, rather than a cloud of microscopic particles, aligning its physical origin squarely with quantum-gravitational microstructure.
How justified are we in considering a holographic origin for dark matter? From a foundational standpoint, there is no conflict in extending the holographic energy density prescription, which has been extensively used for dark energy, to the case of dark matter. The holographic bound itself does not specify what the encoded energy represents and instead it only constrains the total number of degrees of freedom that a region of space can possess without gravitational collapse. Thus, applying the same logic to the clustering sector simply acknowledges that the total energy content of the Universe, including matter, must respect the holographic entropy limit. Extending holography to dark matter is not a violation but a generalization of the same principle across distinct horizon domains. Consequently, HDM preserves all of the principles of cosmological holography while offering a physically coherent, information-theoretic explanation for dark matter that does not the need unknown particles, remnants, or exotic fields. Moreover, in many quantum-gravity frameworks like AdS/CFT [47, 48, 49], causal horizon thermodynamics [50, 51, 52] and covariant entropy bounds [53, 54], the partitioning of bulk energy into “matter-like” and “vacuum-like” components is not fundamental as only the total energy stored in horizon degrees of freedom is fundamental. This is reflected in the covariant entropy bound
$S_{\rm matter}(L)+S_{\rm grav}(L)\leq\frac{A(L)}{4}$ (15)
which does not privilege dark energy over dark matter in any way. Therefore, assigning part of the holographically allowed energy budget to a clustering sector is entirely consistent with these QG principles. Once this horizon-governed energy is projected into the bulk Einstein equations, it naturally behaves as a matter component with scaling determined by the infrared cutoff precisely as realized in the Ricci HDM model we have presented.